# Mr. DETR: Instructive Multi-Route Training for Detection Transformers

### v3

@@ Abstract
Existing methods enhance the training of detection transformers by incorporating an auxiliary one-to-many assignment. In this work, we treat the model as a multi-task framework, simultaneously performing one-to-one and one-to-many predictions. We investigate the roles of each component in the transformer decoder across these two training targets, including self-attention, cross-attention, and feed-forward network. Our empirical results demonstrate that any independent component in the decoder can effectively learn both targets simultaneously, even when other components are shared. This finding leads us to propose a multi-route training mechanism, featuring a primary route for one-to-one prediction and two auxiliary training routes for one-to-many prediction. We enhance the training mechanism with a novel instructive self-attention that dynamically and flexibly guides object queries for one-to-many prediction. The auxiliary routes are removed during inference, ensuring no impact on model architecture or inference cost. We conduct extensive experiments on various baselines, achieving consistent improvements as shown in Fig. 1. Project page: https://visual-ai.github.io/mrdetr
1 Introduction
The end-to-end detection transformer (DETR) [3], along with subsequent studies [34, 41, 53], shine in object detection. Unlike traditional methods, DETR-based detectors eliminate the need for non-maximum-suppression (NMS). They employ the one-to-one assignment strategy (see Fig. 2(a)) for supervision, where each ground-truth box is matched to a single prediction. Compared to the one-to-many assignment methods (see Fig. 2(b)) in conventional object detectors [44, 33, 52] that allow each ground-truth box to be assigned with multiple predictions, the one-to-one assignment can lead to slow convergence [21, 61, 50] of DETRs due to its sparse supervision.
Figure 1: Improvements over different baseline models. The results are reported on the COCO 2017 validation in mean Average Precision (mAP). From left to right: Deformable-DETR++ [70] with 300 queries, H-DETR [21], Deformable-DETR++ [70] with 900 queries, Deformable-DETR++ [70] trained for 24 epochs, DINO [61], Align-DETR [1] and Align-DETR [1] configured with 24 training epochs.
@@ 1 Introduction
Table 1: The AP performance of different variants. Each variant includes a primary route for one-to-one prediction and several auxiliary routes for one-to-many prediction. ‘o2o’: the performance of the primary route. ‘o2m’: the maximum performance with NMS among the auxiliary routes.
@@
In this work, we first build the multi-task framework and empirically investigate the roles of each component in the transformer decoder for one-to-one and one-to-many assignments, including self-attention, cross-attention, and the feed-forward network (FFN). Specifically, we regard the model with auxiliary training by one-to-many assignment as a typical multi-task framework, requiring simultaneous one-to-one and one-to-many predictions. As illustrated in Fig. 3(a), the most straightforward approach is to obtain the prediction results of two tasks with shared components. However, experimental results in Tab. 1(1) indicate that incorporating a one-to-many assignment significantly degrades the performance of the primary one-to-one prediction when all components are shared between two tasks. Consequently, we investigate which component can be independent to enhance the primary route’s performance. We build the multi-task framework with independent self-attention, cross-attention, and FFN for two training targets, as shown in Fig. 3(b-d), respectively. We find that any independent component significantly benefits the primary route of one-to-one prediction, even when other components are shared, achieving 2.1%, 1.6%, and 2.0% improvement, respectively, as demonstrated in Tab. 1(3-5). We also explore the different auxiliary routes with two independent components, which show similar performance to those with a single independent component, but involve more trainable parameters. Therefore, we do not consider auxiliary training routes with two independent components.
@@
This empirical finding motivates us to build a multi-route training mechanism that combines multiple auxiliary training routes, each with an independent component. Specifically, the primary route is used for the one-to-one prediction, while each auxiliary route features a distinct independent component for one-to-many prediction. For instance, in Fig. 3(e), we integrate auxiliary training using independent self-attention and FFN, respectively, with the primary route. We conduct experiments to verify the performance of each combination in Tab. 1. We find that when involving an auxiliary route with independent cross-attention degrades the primary route’s performance, likely due to the slow convergence [50] of independent cross-attention. The combination in Fig. 3(e) outperforms other variants, leading us to adopt this framework as our potential solution.
@@
To further reduce the new trainable parameters in the auxiliary training routes and enhance parameter sharing across all routes, we propose a novel instructive self-attention mechanism, depicted in Fig. 3(f). Our approach includes three training routes, a primary route for one-to-one prediction, and two auxiliary routes with instructive self-attention and independent FFN, respectively. The first auxiliary route shares all parameters with the self-attention in the primary route, but incorporates a learnable token, named instruction token, attached to the input object queries in the self-attention. The object queries, along with the instruction tokens, undergo self-attention, enabling the instruction token to guide the queries for one-to-many prediction. Given that the FFN is implemented as a simple two-layer MLP, we directly employ the independent FFN in the second auxiliary route. During inference, the two auxiliary training routes are discarded, ensuring the model architecture and inference time remain consistent with baseline models.
In summary, we make the following three contributions: First, in the multi-task framework, we empirically demonstrate that any independent component in the decoder can effectively learn one-to-one and one-to-many targets simultaneously, even when other components are shared. Second, building on the above insight, we propose a multi-route training mechanism and enhance it with our proposed instructive self-attention that dynamically and flexibly guides object queries for one-to-many prediction. Third, to verify the effectiveness of our method, we conduct extensive experiments on the COCO 2017 dataset, demonstrating consistent improvements across various baseline models.
2 Related Work
Detection with Transformers. Numerous follow-up works [9, 13, 70] accelerate the DETR [3] convergence by incorporating advanced attention architectures [70, 57, 37, 14, 59, 19, 66] or novel anchor box generation methods [56, 34, 41, 53, 18, 64]. Additionally, some studies propose lightweight detectors [45, 31, 67, 28, 6] or localization-aware objective [62, 35, 1, 19, 43]. In this work, we propose a multi-route training mechanism that enhances the training of DETR models while maintaining the original model architecture during inference.
@@ 2 Related Work
Training Detection Transformers.  Many typical object detectors [44, 52, 63, 33, 16, 68, 12] design various strategies to match each target with multiple predictions, thereby improving the detector’s ability to learn robust representations [21]. In contrast, DETR-like detectors [3, 15, 7, 55, 30, 2] achieve end-to-end object detection through one-to-one matching, where each target is assigned to a single prediction, considering both localization and classification costs. Nonetheless, DETR-like detectors with one-to-one matching frequently encounter slow convergence issues [21, 41, 50]. To accelerate the training of vanilla DETR [3], several studies incorporate auxiliary training methods. For instance, H-DETR [21] leverages Hungarian Matching [25] to construct one-to-many matching by duplicating targets. DN-DETR [27] and DINO [61] design multiple groups of denoising queries to accelerate the training of detection transformers. Likewise, Group-DETR [5] utilizes a set of learnable queries as the auxiliary input for the transformer decoder. DQ-DETR [8] establishes the primary query group by dynamically fusing auxiliary queries. StageInteractor [51] introduces one-to-many matching by combining the label assignment results of different decoder layers. SQR-DETR [4] constructs auxiliary training by reusing object queries from previous decoder layers. Co-DETR [71] adopts multiple query groups with different matching strategies to offer various supervision signals. DAC-DETR [20] develops a parallel decoder to learn a one-to-many assignment by eliminating self-attention. MS-DETR [65] proposes to supervise cross-attention output using one-to-many matching, while self-attention is supervised by one-to-one matching.
@@
Figure 4: Illustration of our proposed multi-route training method. It includes three training routes: Route-1, Route-2, and Route-3. All three routes share the same object queries and detection heads for classification and regression. Route-2 serves as the primary route for one-to-one prediction, identical to the baseline models. Route-1 shares self-attention and cross-attention but uses an independent feed-forward network (o2m FFN) for one-to-many prediction. Route-3, sharing all components with the primary route, introduces a novel instructive self-attention, implemented by adding a learnable instruction token to the object queries to guide them and the subsequent network for one-to-many prediction. During inference, the auxiliary routes, Route-1 and Route-3, are discarded.
3 Method
3.1 Preliminary
@@
One-to-one Training Objective.
Utilizing a one-to-one training objective [3], DETRs achieve end-to-end detection without the need for non-maximum-suppression (NMS). Specifically, let $\mathbf{\bar{B}}=\{\bar{b}_{0},\bar{b}_{1},...,\bar{b}_{t}\}$ and $\mathbf{\bar{S}}=\{\bar{s_{0}},\bar{s_{1}},...,\bar{s_{t}}\}$ represent the ground-truth boxes and corresponding classes. The matching cost between all possible prediction and ground-truth pairs is derived by considering both the classification cost and box costs. Optimal matches are determined using bipartite matching [3, 25], denoted as $\sigma$. The one-to-one training objective is expressed as:
@@
One-to-many Training Objective.
Traditional object detectors [44, 33, 52] typically use a one-to-many assignment strategy, assigning each ground-truth box to multiple predictions based on specific criteria, followed by NMS to eliminate duplicated predictions. In our work, we apply a straightforward one-to-many assignment strategy [42] as in [65, 20], which considers localization quality and classification confidence, making it suitable for DETR-like detectors. Specifically, the matching score $M_{ij}$ between the predictions $(s_{i},b_{i})$ and the ground truth $(\bar{s_{t}},\bar{b_{t}})$ is defined as:
@@
3.2 Multi-route Training
We aim to introduce the one-to-many assignment as an additional training strategy to enhance detection transformers. First, we treat the detector with auxiliary one-to-many prediction as a multi-task framework, which simultaneously achieves both one-to-one and one-to-many predictions. As discussed in Sec. 1, we empirically investigate the roles of each component in the transformer decoder within this framework and find: (i) incorporating a one-to-many assignment significantly degrades the performance of the primary one-to-one prediction when all components are shared between two tasks (see Tab. 1(2)). We anticipate that it results from the interference between two tasks. For instance, a predicted box may be assigned as a positive prediction in the one-to-many assignment, while being assigned as a negative prediction in the one-to-one assignment. (ii) any independent component in the decoder significantly benefits the primary one-to-one prediction route, even when other components are shared (see Fig. 3 (b - d) & Tab. 1 (3-5)). This observation indicates that any independent component is capable of effectively mastering one-to-one and one-to-many training goals, thereby resolving the conflict between these two tasks. This is expected since shared components can extract common clues for two tasks, while independent components further distinguish different tasks; (iii) an auxiliary training route with two independent components does not outperform the route with only one independent component (see Tab. 1(6-8)); and (iv) combining the auxiliary route with independent self-attention and independent FFN achieves the highest performance among variants that combine different auxiliary training routes (see Tab. 1(10)).

### v4

# Mr. DETR++: Instructive Multi-Route Training for Detection Transformers with Mixture-of-Experts

@@ Abstract
Existing methods enhance the training of detection transformers by incorporating an auxiliary one-to-many assignment. In this work, we treat the model as a multi-task framework, simultaneously performing one-to-one and one-to-many predictions. We investigate the roles of each component in the transformer decoder across these two training targets, including self-attention, cross-attention, and feed-forward network. Our empirical results demonstrate that any independent component in the decoder can effectively learn both targets simultaneously, even when other components are shared. This finding leads us to propose a multi-route training mechanism, featuring a primary route for one-to-one prediction and two auxiliary training routes for one-to-many prediction. We propose a novel instructive self-attention mechanism, integrated into the first auxiliary route, which dynamically and flexibly guides object queries for one-to-many prediction. For the second auxiliary route, we introduce a route-aware Mixture-of-Experts (MoE) to facilitate knowledge sharing while mitigating potential conflicts between routes. Additionally, we apply an MoE to low-scale features in the encoder, optimizing the balance between efficiency and effectiveness. The auxiliary routes are discarded during inference. We conduct extensive experiments across various object detection baselines, achieving consistent improvements as demonstrated in Fig. 1. Our method is highly flexible and can be readily adapted to other tasks. To demonstrate its versatility, we conduct experiments on both instance segmentation and panoptic segmentation, further validating its effectiveness. Project page: https://visual-ai.github.io/mrdetr/
Index Terms:
Object Detection, detection transformers, multi-route training, instructive self-attention, mixture-of-experts.
I Introduction
The end-to-end detection transformer (DETR) [1], along with subsequent studies [2, 3, 4], shine in object detection due to its simplicity and effectiveness. Unlike traditional methods, DETR-based detectors eliminate the need for post-processing steps like non-maximum-suppression (NMS). They employ the one-to-one assignment strategy (see Fig. 2(a)) for supervision, where each ground-truth box is matched to a single prediction. Compared to the one-to-many assignment methods (see Fig. 2(b)) in conventional object detectors [5, 6, 7] that allow each ground-truth box to be assigned with multiple predictions, the one-to-one assignment can lead to slow convergence [8, 9, 10] of DETRs due to its sparse supervision.
Figure 1: Improvements over different baseline models. The results are reported on the COCO 2017 validation in mean Average Precision (mAP). From left to right: Deformable-DETR++ [11] with 300 queries, Deformable-DETR++ [11] with 900 queries, Deformable-DETR++ [11] trained for 24 epochs, DINO [9], DINO trained for 24 epochs, Align-DETR [12] and Align-DETR [12] configured with 24 training epochs.
@@ I Introduction
TABLE I: The AP performance of different variants. Each variant includes a primary route for one-to-one prediction and several auxiliary routes for one-to-many prediction. ‘o2o’: the performance of the primary route. ‘o2m’: the maximum performance with NMS among the auxiliary routes.
@@
In this work, we first build the multi-task framework and empirically investigate the roles of each component in the transformer decoder for one-to-one and one-to-many assignments, including self-attention, cross-attention, and the feed-forward network (FFN). Specifically, we regard the model with auxiliary training by one-to-many assignment as a typical multi-task framework, requiring simultaneous one-to-one and one-to-many predictions. As illustrated in Fig. 3(a), the most straightforward approach is to obtain the prediction results of two tasks with shared components. However, experimental results in Tab. I(1) indicate that incorporating a one-to-many assignment significantly degrades the performance of the primary one-to-one prediction when all components are shared between two tasks. Consequently, we investigate which component can be independent to enhance the primary route’s performance. We build the multi-task framework with independent self-attention, cross-attention, and FFN for two training targets, as shown in Fig. 3(b-d), respectively. We find that any independent component significantly benefits the primary route of one-to-one prediction, even when other components are shared, achieving 2.1%, 1.6%, and 2.0% improvement, respectively, as demonstrated in Tab. I(3-5). We also explore the different auxiliary routes with two independent components, which show similar performance to those with a single independent component, but involve more trainable parameters. Therefore, we do not consider auxiliary training routes with two independent components.
@@
This empirical finding motivates us to build a multi-route training mechanism that combines multiple auxiliary training routes, each with an independent component. Specifically, the primary route is used for the one-to-one prediction, while each auxiliary route features a distinct independent component for one-to-many prediction. For instance, in Fig. 3(e), we integrate auxiliary training using independent self-attention and FFN, respectively, with the primary route. We conduct experiments to verify the performance of each combination in Tab. I. We find that when involving an auxiliary route with independent cross-attention degrades the primary route’s performance, likely due to the slow convergence [10] of independent cross-attention. The combination in Fig. 3(e) outperforms other variants, leading us to adopt this framework as our potential solution.
@@
To reduce additional trainable parameters in the auxiliary training routes and promote parameter sharing across all routes, we propose a novel instructive self-attention mechanism, depicted in Fig. 3(f) named Mr. DETR. It includes three training routes: a primary route for one-to-one prediction and two auxiliary routes, incorporating instructive self-attention and an independent FFN, respectively. The first auxiliary route shares all parameters with the self-attention in the primary route, but incorporates a learnable token, named instruction token, attached to the input object queries in the self-attention. The object queries, along with the instruction token, undergo self-attention, enabling the instruction token to guide the queries for one-to-many prediction. We directly employ the independent FFN in the second auxiliary route. During inference, the two auxiliary training routes are discarded, ensuring the model architecture and inference time remain consistent with baseline models.
To share knowledge between two independent FFNs designed as distinct experts for different prediction targets, we further propose a route-aware mixture-of-experts (MoE) framework, with which Mr. DETR evolves into Mr. DETR++, by replacing the two independent FFNs with the proposed route-aware MoE (see Fig. 3(g)). Mr. DETR++ incorporates three routes: a primary route for one-to-one prediction and two auxiliary routes for one-to-many prediction. In this architecture, all routes share a group of experts, with the primary route and one auxiliary route (employing instructive self-attention) sharing the same gating mechanism, while another auxiliary route uses independent gating to mitigate conflicts between routes. This design enables the route-aware MoE to share knowledge across experts while avoiding conflicts between routes. To extend the MoE framework to the transformer encoder, we introduce a scale-aware MoE strategy to address the computational cost of processing long sequences formed by flattened multi-scale image features. This strategy shares an expert across all tokens and disables gated experts for higher scales, significantly reducing computational overhead. Additionally, we propose a localization-aware calibrating strategy that, beyond the vanilla classification score, learns an IoU-related score during training, inspired by VarifocalNet [18], to jointly assess classification and localization quality. Unlike Stable-DINO [19] and Rank-DINO [20], this score is excluded from the label assignment process, enabling it to serve as a simple plug-in module. During inference, the localization-aware score calibrates the classification score, enhancing prediction accuracy.
In summary, we make the following main contributions:
  • We demonstrate empirically that, within a multi-task framework, any independent component in the decoder can effectively learn both one-to-one and one-to-many prediction targets simultaneously, even when other components are shared.
  • Leveraging this insight, we propose a multi-route training mechanism enhanced by a novel instructive self-attention mechanism that dynamically guides object queries for one-to-many prediction, culminating in the development of Mr. DETR.
  • We introduce a route-aware mixture-of-experts (MoE) framework to enable knowledge sharing between the two independent feed-forward networks (FFNs) in Mr. DETR. By further incorporating a scale-aware MoE in the transformer encoder, we develop Mr. DETR++.
  • We investigate the intrinsic mechanisms of our proposed multi-route training framework through probing experiments. Our findings reveal that auxiliary training routes mitigate conflicts with the primary route by selectively disrupting information necessary for one-to-one prediction within object queries.
  • We validate the effectiveness of our approach through extensive experiments on multiple benchmarks, including COCO 2017, Objects365, and NuImages, achieving consistent improvements over various baseline models. Additionally, we extend our method to diverse tasks, such as instance segmentation and panoptic segmentation.
A preliminary version of this work has been presented in [21], which investigates the role of transformer decoder components in a multi-task framework and introduces a multi-route training mechanism with a novel instructive self-attention mechanism. In this work, we further propose a route-aware MoE framework to enable knowledge sharing while mitigating conflicts between routes for one-to-one and one-to-many predictions. Additionally, we introduce a scale-aware MoE in the transformer encoder to enhance feature extraction capabilities. By incorporating a localization-aware score, we develop a simple yet effective score calibration method. To demonstrate the effectiveness of our approach, we provide extensive experimental results on additional benchmarks, including Objects365 [22] and NuImages [23], and extend our method to the panoptic segmentation task. Furthermore, we conduct probing experiments to thoroughly investigate the intrinsic mechanisms of our proposed multi-route training framework.
II Related Work
II-A Object Detection with Transformers
Unlike traditional CNN-based object detectors [5, 24, 7, 25], DETR [1] achieves end-to-end object detection without requiring any post-processing, leveraging a query-based transformer and set prediction mechanism. To enhance performance and accelerate the training of DETR [1], numerous follow-up works [26, 27, 11] introduce advanced attention architectures. Specifically, Deformable-DETR [11] introduces deformable attention to enable efficient multi-scale attention. Cascade-DETR [28] improves the performance of deformable attention by utilizing predicted boxes. EASE-DETR [29] enhances one-to-one prediction by modulating self-attention with the location relationship of queries, while Relation-DETR [30] achieves similar improvements by incorporating location relationships across different decoder layers. SAP-DETR [31] proposes a modulated attention architecture augmented with additional salient points. Furthermore, SAM-DETR [32], HPR [33], and MI-DETR [34] integrate RoI-Align [5] or multiple cross-attention mechanisms to enhance object queries in the transformer decoder. Another line of research focuses on designing novel anchor box generation mechanisms [35, 2, 3, 4, 36, 37]. For example, DAB-DETR [2], Conditional-DETR [3], and Anchor-DETR [4] improve the transformer decoder’s capability by associating anchor boxes with object queries. Efficient-DETR [35], Salience-DETR [36], and DDQ-DETR [37] further enhance detection by imposing a localization prior to generate stronger object proposals. Additionally, lightweight DETR models [38, 39, 40, 41, 42] and localization-aware training objectives [43, 19, 12, 30, 20] have been proposed to improve efficiency and accuracy. In our work, we propose an instructive multi-route training mechanism that enhances the training of various DETR-like models.
II-B Training Detection Transformers
@@
Many typical object detectors [5, 7, 44, 6, 45, 24, 46] design various strategies to match each target with multiple predictions, thereby improving the detector’s ability to learn robust representations [8]. In contrast, DETR-like detectors [1, 47, 48, 49, 50, 51, 52] achieve end-to-end object detection through one-to-one matching, where each target is assigned to a single prediction, considering both localization and classification costs. Nonetheless, DETR-like detectors with one-to-one matching frequently encounter slow convergence issues [8, 3, 10]. To accelerate the training of vanilla DETR [1], several studies incorporate auxiliary training methods. For instance, H-DETR [8] leverages Hungarian Matching [53] to construct one-to-many matching by duplicating targets. DN-DETR [15] and DINO [9] design multiple groups of denoising queries to accelerate the training of detection transformers. Likewise, Group-DETR [16] utilizes a set of learnable queries as the auxiliary input for the transformer decoder. DQ-DETR [54] establishes the primary query group by dynamically fusing auxiliary queries. StageInteractor [55] introduces one-to-many matching by combining the label assignment results of different decoder layers. SQR-DETR [56] constructs auxiliary training by reusing object queries from previous decoder layers. Co-DETR [57] adopts multiple query groups with different matching strategies to offer various supervision signals. DAC-DETR [13] develops a parallel decoder to learn a one-to-many assignment by eliminating self-attention. MS-DETR [14] proposes to supervise cross-attention output using one-to-many matching, while self-attention is supervised by one-to-one matching. Additionally, several works [58, 59, 60] use knowledge distillation [61, 62] to enhance the training of DETR models. Different from previous arts, we discuss the roles of each component in the transformer decoder in the multi-task framework, which is required to achieve one-to-one and one-to-many prediction. We further propose an instructive multi-route training mechanism for the utilization of auxiliary training.
II-C Mixture-of-Experts
Mixture-of-Experts (MoE) [63] aims to dynamically combine the knowledge of multiple experts while maintaining low computational cost during inference. This technique has been widely applied across various domains to scale up models, such as natural language processing [64, 65, 66] and computer vision [67, 68, 69, 70, 71]. In the context of multi-task learning, M3ViT [72] and Mod-Squad [73] introduce MoE-based models to effectively handle multiple tasks. AdaMV-MoE [74] enhances task specificity while maintaining efficiency through the use of partially shared experts. Furthermore, [75] proposes constructing experts with LoRA [76] to address multiple dense prediction tasks.
@@ II-C Mixture-of-Experts
Figure 4: Our proposed multi-route training method in Mr. DETR. It includes three training routes: Route-1, Route-2, and Route-3. All three routes share the same object queries and detection heads for classification and regression. Route-2 serves as the primary route for one-to-one prediction, identical to the baseline models. Route-1 shares self-attention and cross-attention but uses an independent feed-forward network (o2m FFN) for one-to-many prediction. Route-3, sharing all components with the primary route, introduces a novel instructive self-attention, implemented by adding a learnable instruction token to the object queries to guide them and the subsequent network for one-to-many prediction. During inference, the auxiliary routes, Route-1 and Route-3, are discarded.
III Method
III-A Preliminary
@@
One-to-one Training Objective. Utilizing a one-to-one training objective [1], DETRs achieve end-to-end detection without the need for non-maximum-suppression (NMS). Specifically, let $\mathbf{\bar{B}}=\{\bar{b}_{0},\bar{b}_{1},...,\bar{b}_{t}\}$ and $\mathbf{\bar{S}}=\{\bar{s_{0}},\bar{s_{1}},...,\bar{s_{t}}\}$ represent the ground-truth boxes and corresponding classes. The matching cost between all possible prediction and ground-truth pairs is derived by considering both the classification cost and box costs. Optimal matches are determined using bipartite matching [1, 53], denoted as $\sigma$. The one-to-one training objective is expressed as:
@@
One-to-many Training Objective. Traditional object detectors [5, 6, 7] typically use a one-to-many assignment strategy, assigning each ground truth box to multiple predictions based on specific criteria, followed by NMS to eliminate duplicated predictions. In our work, we apply a straightforward one-to-many assignment strategy [17] as in [14, 13], which considers localization quality and classification confidence, making it suitable for DETR-like detectors. Specifically, the matching score $M_{ij}$ between the predictions $(s_{i},b_{i})$ and the ground truth $(\bar{s_{t}},\bar{b_{t}})$ is defined as:
@@
III-B Multi-route Training
We aim to introduce the one-to-many assignment as an additional training strategy to enhance detection transformers. First, we treat the detector with auxiliary one-to-many prediction as a multi-task framework, which simultaneously achieves both one-to-one and one-to-many predictions. As discussed in Sec. I, we empirically investigate the roles of each component in the transformer decoder within this framework and find: (i) incorporating a one-to-many assignment significantly degrades the performance of the primary one-to-one prediction when all components are shared between two tasks (see Tab. I(2)). We anticipate that it results from the interference between two tasks. For instance, a predicted box may be assigned as a positive prediction in the one-to-many assignment, while being assigned as a negative prediction in the one-to-one assignment. (ii) any independent component in the decoder significantly benefits the primary one-to-one prediction route, even when other components are shared (see Fig. 3 (b - d) & Tab. I (3-5)). This observation indicates that any independent component is capable of effectively mastering one-to-one and one-to-many training goals, thereby resolving the conflict between these two tasks. This is expected since shared components can extract common clues for two tasks, while independent components further distinguish different tasks; (iii) an auxiliary training route with two independent components does not outperform the route with only one independent component (see Tab. I(6-8)); and (iv) combining the auxiliary route with independent self-attention and independent FFN achieves the highest performance among variants that combine different auxiliary training routes (see Tab. I(10)).